OpenStudy (anonymous):

Graph the equation. y = 4x – 3

OpenStudy (anonymous):

put 2-3 different values of x and get the corresponding values of y. then plot them and connect the dots n u hv yr graph.... e.g. if u put x=0, then y=4*0-3 so y= -3 so point is (0, -3) if u put x = 2, then y=4*2 - 3 so y = 8-3=5 so point is (2, 5)

OpenStudy (anonymous):

Another option is to remember that your y-intercept here is simply -3. To find the x-intercept set y = 0, and solve for x. Then you can simply connect the two points.